OpenStudy (anonymous):

can someone explain to me how to do this : an = ½ an-1 , where a1 = 20

OpenStudy (tkhunny):

\(a_{n} = f(a_{n-1})\) MEANS the following sequence... \(a_{2} = f(a_{1})\) \(a_{3} = f(a_{2})\) \(a_{4} = f(a_{3})\) \(a_{5} = f(a_{4})\) \(a_{6} = f(a_{5})\) ... You are given \(a_{1}\). Apply the formula (multiply by 1/2) to obtain \(a_{2}\) You now have \(a_{2}\). Apply the formula (multiply by 1/2) to obtain \(a_{3}\) and so forth...

OpenStudy (anonymous):

soo all i have to do is multiply 20 by 1/2 ?

OpenStudy (tkhunny):

...and that will be \(a_{2}\). Do it again to find \(a_{3}\).

OpenStudy (tkhunny):

Oftentimes, the formula is more complicated.

OpenStudy (anonymous):

wait a sec

OpenStudy (anonymous):

okay so I got 10 for a2 is that right?

OpenStudy (tkhunny):

\(a_{1} = 20\) \(a_{2} = (1/2)a_{1}\) \(a_{2} = (1/2)\cdot 20 = 10\) \(a_{2} = 10\) I think you have it. Why do you doubt?

OpenStudy (anonymous):

okay, im still confused is that the answer of -->an = ½ an-1<--

OpenStudy (tkhunny):

You missed an important part of the discussion. Repeated here for your convenience. "MEANS the following sequence..." That notation suggests a sequence of events. Another, perhaps more clear, way to write it might be, \(a_{NextInSequence} = (1/2)⋅(a_{PreviousInSequence})\) You need only a place to start, such as a1=20, in order to define the entire SEQUENCE. To generate the ENTIRE sequence, you must apply the formula infinitely many times.

OpenStudy (jdoe0001):

\(\bf \large{a_{\color{red}{ n}}=\cfrac{1}{2}a_{{\color{red}{ n}}-1}\qquad \qquad a_1=20\\ \quad \\ a_{\color{red}{ 2}}=\cfrac{1}{2}a_{{\color{red}{ 2}}-1}\implies a_{\color{red}{ 2}}=\cfrac{1}{2}a_{{\color{red}{ 1}}}\qquad a_{\color{red}{ 3}}=\cfrac{1}{2}a_{{\color{red}{ 3}}-1}\implies \cfrac{1}{2}a_{{\color{red}{ 2}}}\\ \quad \\ a_{\color{red}{ 7}}=\cfrac{1}{2}a_{{\color{red}{ 7}}-1}\implies \cfrac{1}{2}a_{{\color{red}{ 6}}}\qquad a_{\color{red}{ 9}}=\cfrac{1}{2}a_{{\color{red}{ 9}}-1}\implies \cfrac{1}{2}a_{{\color{red}{ 8}}}\quad ...}\)

OpenStudy (anonymous):

ohh okay i get it now thank you yall